MR. BARNES STOCKED HIS GARDEN SUPPLY STORE WITH 2 TONS OF SAND IN 50-POUND BAGS FOR $138 PER TON. IF HE SELLS THE ENTIRE ORDER OF BAGS OF SAND FOR $5.95 PER BAG, WHAT WILL BE HIS PROFIT?
Answer by Theo(13342)

he bought 2 tons of sand at 138 per ton.
his cost was 2 * 138 = 276.

he puts the 2 tons of sand in 50 pound bags.
1 ton is equal to 2000 pounds.
2 tons is equal to 4000 pounds.
number of bags he sells is 4000 / 50 = 80.

80 bags at 5.95 per bag = 476 dollars.

his cost was 276.
his revenue was 476.
his profit is 46 - 276 = 200 dollars.
that's your solution.